OpenStudy (anonymous):

Describe the graphs of the functions f(x) = 2^x – 3 and g(x) = –2^x – 3. Compare and contrast the domain and range of f(x) and g(x) Describe the graphs of the functions f(x) = 2^x – 3 and g(x) = –2^x – 3. Compare and contrast the domain and range of f(x) and g(x) @Mathematics

OpenStudy (turingtest):

f(x) = 2^x – 3 y-intercept=-2 horizontal asymptote: y=-3 Domain all reals, Range y ≥-3

OpenStudy (turingtest):

g(x) = –2^x – 3 y-int=-2 horizontal asymptote=-3 Domain ALL REALS Range y≤-3

OpenStudy (turingtest):

f(x) increases as x increases g(x) decreases as x increases

OpenStudy (anonymous):

Thank you :) how did you find the range?

OpenStudy (turingtest):

Imagine if \[x \to -\infty\]then \[2^x \to 0\]which leave the equations as f(x)=-3 and g(x)=-3. so they cannot go past -3.

OpenStudy (turingtest):

actually I made a slight mistake since x cannot equal infinity, f(x) and g(x) cannot EQUAL 3, they can only approach it asymptotically, so the ranges should be f(x): y>-3 and g(x) y<-3

OpenStudy (anonymous):

Oh ok, so 3 is like the discontinuity?

OpenStudy (turingtest):

Well, a discontinuity is a point where the graph has a "hole" or a "jump". This is not the case here, since the function is continuous. The point is that making x more and more negative only APPROACHES 3, but never equals it. I presume you have not studied limits yet, but the idea will become more clear when you do.

OpenStudy (anonymous):

Mhm, thanks for your help, and sorry im not to good with ranges in graphs haha.

OpenStudy (anonymous):

So if the line goes on the number and goes in a certain direction, that is how you find the range?

OpenStudy (turingtest):

Yeah many ranges can only be determined through Calculus, but some things to watch for are square roots and exponential functions, like \[y=\sqrt x\] has a range of \[[0,\infty)\]because if x<0 y is imaginary (In this case the domain is the same as the range. Think about why.) Or\[y=a^x\]has a range of \[(0,\infty]\]notice that 0 is not included in the range because in order to get \[y=0\] we must have\[ x=-\infty\] which we cannot do (plugging in infinity as a value is generally a no-no in basic mathematics)

OpenStudy (turingtest):

I have no idea what you meant by your last question...

OpenStudy (turingtest):

The range is the set of all possible values a function can take on. I'm not sure I can describe it much beyond that.

OpenStudy (anonymous):

Kk well thank you for your help :)

OpenStudy (anonymous):

Oh I totally get it now, I had to graph it -__- lol thank you so much